OpenStudy (anonymous):

simplify radical 75 - radical 147-5 radical 8 + 3 radical 48

OpenStudy (anonymous):

Simplify \[\sqrt{75}-\sqrt{147}- 5\sqrt{8} +3\sqrt{48}\]

OpenStudy (anonymous):

possible anwsers

OpenStudy (anonymous):

\[-2\sqrt{35} or \sqrt{5} or -8+5\sqrt{6} or -4\sqrt{2}+5\sqrt{3} \]

OpenStudy (anonymous):

\[=\sqrt{25}\sqrt{3}-\sqrt{49}\sqrt{3}-5\sqrt{4}\sqrt{2}+3\sqrt{16}\sqrt{3}\] \[=5\sqrt{3}-7\sqrt{3}-10\sqrt{2}+12\sqrt{3}\] \[=10\sqrt{3}-10\sqrt{2}\] Does this help?

OpenStudy (anonymous):

yea a little thanks alot i can take it from here

OpenStudy (anonymous):

No problem :)